Hi,
I tried to do this problem, and I would just like to check with you.
The problem is:
P=2L+2W solve for L
so i subtracted 2W from both sides
and got: P-2W=2L
then i divided by 2 and got
L=P-2W/2
am i correct? or is there something else that needs to be done?
thanks in advance

Found 2 solutions by stanbon, checkley77:
Answer by stanbon(75887)

I tried to do this problem, and I would just like to check with you.
The problem is:
P=2L+2W solve for L
so i subtracted 2W from both sides
and got: P-2W=2L
then i divided by 2 and got
L=P-2W/2
am i correct? or is there something else that needs to be done?
------------------
Since the numerator has two terms, you need parenthses there.
L = (P-2W)/2
==================
Cheers,
Stan H.

Answer by checkley77(12844)

P=2L+2W solve for L
so i subtracted 2W from both sides
and got: P-2W=2L
then i divided by 2 and got
L=(P-2W)/2 YOU FORGOT THE () AROUND THE NUMERATOR.
am i correct? or is there something else that needs to be done?
thanks in advance